EDITORIAL

Terry Goodman

As is our custom, we want to acknowledge those who have served as referees for MJMS for the past year. Their contribution to the journal is so very important and we appreciate the time and energy they give to their task of reviewing submitted manuscripts.

List of Referees

Mangho Ahuja, Cape Girardeau, Missouri Elvan Akin, University of Missouri–Rolla Hadi Y. Alkahby, Dillard University Jamuna P. Ambasht, Benedict College Mohammad K. Azarian, University of Evansville Dale Bachman, Central Missouri State University Seung–Jin Bang, Seoul, Korea Miron Bekker, University of Missouri–Rolla James Benedict, August State College Gerald Bergum, Brookings, South Dakota John E. Bernard, University of Texas–Pan American David A. Brown, Ithaca College Reginald Brigham, University of Missouri–Rolla Jim Bruening, Southeast Missouri State University Larry Cammack, Central Missouri State University G. Alan Cannon, Southeastern Louisiana University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Steve Clark, University of Missouri–Rolla Jennifer Fowler Daniel, Lamar University Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Thomas Dence, Ashland University Albert Dixon, College of the Ozarks Clayton Dodge, University of Maine

List of Referees (cont.)

Roman Dwilewicz, University of Missouri–Rolla John J. Edgell, Jr., Southwest Texas State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, San Antonio, Texas Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, University of Central Arkansas Lou Grimm, University of Missouri–Rolla Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State University Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Dean Hoffman, Auburn University Joe Howard, New Mexico Highlands University Timothy P. Keller, George Mason University Robert E. Kennedy, Pleasant Hill, Missouri Sava Krstic, Oregon Health & Science University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Michael Laidacker, Lamar University Mulatu Lemma, Savannah State University Edward S. Letzter, Temple University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State University Roger Marty, Cleveland State University Alec Matheson, Lamar University Katherine Mawhinney, Appalachian State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Ilene Morgan, University of Missouri–Rolla

List of Referees (cont.)

Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Kent M. Neuerburg, Southeastern Louisiana University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Bill Read, Lamar University Don Redmond, Southern Illinois University Les Reid, Southwest Missouri State University Jeffrey A. Rosoff, Gustavus Adolphus College Paul Runnion, University of Missouri–Rolla Prasanna Sahoo, University of Louisville Suman Sanyal, University of Missouri–Rolla Jau–Shyong Shiue, University of Nevada–Las Vegas Donald P. Skow, University of Texas–Pan American James Sly, University of Missouri–Rolla Neil Oliver Smith, Augusta State College Allen Stenger, Alamogordo, New Mexico Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Howard Warth, University of Missouri–Rolla Joseph Wiener, University of Texas–Pan American Bobby Neal Winters, Pittsburg State University Christina Welch, University of Missouri–Rolla Fred Worth, Henderson State University